4. The Earth is not neutral. It has a charge of roughly -6.76 x 105 C, which is roughly uniformly distributed over the Earth. (a) How large is the force that the Eartlh exerts on an electron near its surface? Is this force attractive or repulsive? Compare this force with the gravitational force that the Earth exerts on the electron. (b) Suppose we wanted to levitate a paper clip using the electric force that the Earth exerts on it. The paper clip has a mass of 1o g. How much charge must be given to the paper clip (and what sign) for the paper clip to be able to float motionless in the air. (c) How many electrons need to be added/removed from the paper clip to achieve the outcome of part 4b? Express this as a fraction of the total number of electrons in the paper clip. Does it seem like this would be a practically achievable thing to do?
Answer #1

(a)

Force exerted by earth on electron,

F = kq1q2 / r^2

where q1 = charge on earth

q2 = charge on electron

r = radius of earth (distance b/w electron and earth surface)

so F = 9*109*6.76*105*1.6*10-19 / (6.37*106)2

F = 2.39*10-17 N

It is repulsive force as both earth and electron have same charge (negative).

Gravitational force,

Fg = GmM / r^2

Fg = 6.67*10-11*9.1*10-31*5.97*1024 / (6.37*106)2

F = 8.92*10-30 N

Force exerted by earth on electron is much greater than gravitational force.

(b)

Balance the forces on paper clip,

kq1q2 / r2 = mg

9*109*6.76*105*q /  (6.37*106)2 = 1*10-3*9.8

q = 1.61*10-18 C

Sign is negative (to be repaled by earth).

(c)

Number of electrons in paper clip,

n = q / e =  1.61*10-18 / 1.6*10-19

n = 10